1. By what mass (in grams) will a chromium cathode increase when it is coupled to a magnesium half-cell in which the magnesium anode loses 1.53 grams? (assuming chromium ions in solution are Cr3+).


1
Expert's answer
2022-06-14T02:56:03-0400

Equivalent (Mg) = M / valence = 24/2 = 12

Equivalent (Cr) = M / valence = 54/3 = 18


the substances at the cathode and anode are present in equivalent amounts.

12 g — 18 g

1,53 g — X

X = 1,53 * 18 / 12 = 2,295 g


Answer: 2,295 g